OpenStudy (anonymous):

square root 1+tan^2x/1-sin^2x all is under the square root. i did sec^2theta/cos^2theta whats next?

OpenStudy (anonymous):

sec^4 2x

OpenStudy (anonymous):

you just need this trigonometric function to be solved or do you have anything to be proved?

OpenStudy (anonymous):

find what it is equivalent to

OpenStudy (anonymous):

then that's the answer! :)

OpenStudy (anonymous):

\[\sqrt{1+\tan^2x}\div 1-\sin^2x\]

OpenStudy (anonymous):

both of these r under the square root

OpenStudy (anonymous):

ahh sorry i didn't notice that!then the answer is sec^2 x

OpenStudy (anonymous):

how did u get that?

OpenStudy (anonymous):

1+tan^2 x= sec^2 x (that's trigonometry theory) sin^x + cos^x=1 (that's trigonometry theory) therefore [1-sin ^2 x= cos ^2 x] so √1+tan2x ÷ √1−sin2x = √sec^2 x ÷ √cos ^2 x =sec x ÷ cos x =sec ^2 x (because 1÷cos x =sec x)

OpenStudy (anonymous):

oh i c i understand up to the 4th line and i got the answer on my own but up to the fourth line .wcan u explain line 5 and 6 for me .how did u cancel it and set up the fraction,.

OpenStudy (anonymous):

=sec x ÷ cos x = sec x * 1/ cos x = sec x * sec x ( because 1/cos x = sec x) =sec ^2 x

OpenStudy (anonymous):

k thanks so much and thanks for your patience .